OpenStudy (anonymous):

Identify the factors of x2 − 5x − 24. (x + 8)(x − 3) (x − 8)(x + 3) (x + 4)(x − 6) (x − 4)(x + 6)

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

ME HELP

OpenStudy (anonymous):

ok

OpenStudy (haseeb96):

what do you think which one is correct?

OpenStudy (anonymous):

\[x^2-5x-24\] ac=-24

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

can u start from here?

OpenStudy (anonymous):

Wow......

OpenStudy (anonymous):

wait no its A

OpenStudy (haseeb96):

 x^2 - 8x +3x -24 x(x-8) +3(x-8) (x+3)(x-8) is your answer

OpenStudy (anonymous):

cause 8+(-3)=5 and 8*(-3)=-24

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks